Blue Ribbon, Inc. wants to have a weighted average cost of capital of 10 percent. The firm has an after-tax cost of debt of 4 percent and a cost of equity of 12 percent. What debt-equity ratio is needed for the firm to achieve their targeted weighted average cost of capital?The firm face a tax rate of 40%.
Answer #1

Debt- Equity Ratio needed is   0.33

Step-1:Calculation of weight of debt and Equity
Weighted Average cost of capital = (Wd*Kd)+(We*Ke)
0.10 = (w*0.04)+((1-w)*0.12)
0.10 = (0.04w)+(0.12-0.12w)
0.10 = 0.04w+0.12-0.12w
0.10 = 0.12-0.08w
0.08w = 0.02
w = 0.25
so, 1-w = 0.75
Note:
After-tax cost of debt is used in calculation of Weighted Average cost of capital.
Further, sum of weight is always 1. So, if weight of debt is assumed as "w" then weight of equity will be "1-w".
Step-2:Calculation of debt-Equity Ratio
Debt-Equity Ratio = Weight of debt / Weight of Equity
= 0.25 / 0.75
= 0.33